OpenStudy (nincompoop):

tricky or nah? \(\huge \lim_{x \rightarrow 0^+} (\sin x)^{\frac{1}{\ln x}} \)

OpenStudy (abb0t):

Not at all.

OpenStudy (nincompoop):

:)

OpenStudy (kainui):

Did you mean this \[\huge \lim_{x \rightarrow 0^+} \sin (x^{\frac{1}{\ln x}}) \] or this \[\huge \lim_{x \rightarrow 0^+} (\sin x)^{\frac{1}{\ln x}} \]

OpenStudy (nincompoop):

second one hahaha this is for the people who's just taken calc1 not the god level like you, kai

ganeshie8 (ganeshie8):

I wonder if there is a clever way to work this with out using Lhopital

OpenStudy (nincompoop):

without you mean? I do not know without using L'Hospital, but I know using L'Hospital

ganeshie8 (ganeshie8):

If I may use the small angle approximation, \(\sin x\approx x\), and prentend to ignore all technicalities of singularities : \[x^{1/\ln x} = e^{\ln x/\ln x} = e\]

OpenStudy (nincompoop):

oooh that's interesting

ganeshie8 (ganeshie8):

That gives us some number but it raises many serious issues, we won't know if that number is correct or not w/o addressing them...

OpenStudy (nincompoop):

here's what I thought: start out by \(\large y(x) = (sin~x)^{1/ln~x} \rightarrow ln~y = (sin~x)^{1/ln~x} \rightarrow \color{blue}{ln~y = \frac{ln~(sin~x)}{ln~x}}\) then work the limit of \(ln~y \) giving indeterminate form -inf/-inf

OpenStudy (nincompoop):

gives us e

ganeshie8 (ganeshie8):

that is a less unorthodox... and nobody should have any issues with that :)

OpenStudy (nincompoop):

typo \(\large y(x) = (sin~x)^{1/ln~x} \rightarrow ln~y = ln (sin~x)^{1/ln~x} \rightarrow \color{blue}{ln~y = \frac{ln~(sin~x)}{ln~x}} \)

ganeshie8 (ganeshie8):

we may directly use this property of continuous functions : \[\lim f(g(x)) = f(\lim g(x))\]

ganeshie8 (ganeshie8):

\[\lim e^{g(x)} = e^{\lim g(x)}\]

ganeshie8 (ganeshie8):

\[\lim (\sin x)^{1/\ln x}\\~\\ =\lim e^{\ln (\sin x)^{1/\ln x}}\\~\\ =\lim e^{\ln (\sin x)/\ln x}\\~\\ = e^{\lim \ln(\sin x)/\ln x} \]

ganeshie8 (ganeshie8):

thats not much different than what you're doing, but it certanly looks better because we're not introducing new variables and it allows you to go in single direction...

OpenStudy (nincompoop):

I want to medal you again

ganeshie8 (ganeshie8):

Haha all that is possible because of this \[x = e^{\ln x}\]

ganeshie8 (ganeshie8):

assuming x > 0

OpenStudy (solomonzelman):

Continuing where ganeshie8 left off. \(\color{#000000 }{ \displaystyle {\lim_{x \to 0}~\ln(\sin (x))/\ln(x)} }\) (-∞)/(-∞), so L'H'S is applied. \(\color{#000000 }{ \displaystyle {\lim_{x \to 0}~\frac{\frac{\cos(x)}{\sin(x)}}{\frac{1}{x}}} }\) 0/0, so L'H'S is applied again \(\color{#000000 }{ \displaystyle {\lim_{x \to 0}~\frac{\frac{-1}{\sin^2(x)}}{\frac{1}{-x^2}}}={\lim_{x \to 0}~\frac{x^2}{\sin^2x} }=\left[{\lim_{x \to 0}~\frac{x}{\sin x} }\right]^2 }\) \(\color{#000000 }{ \displaystyle }\)

OpenStudy (solomonzelman):

and e^0 ... blah blah blah

OpenStudy (anonymous):

lol why